OpenStudy (anonymous):

Can someone actually show their work so I can see what I did wrong. (1/6x^2) = (1/2x) + (7/6x^2)

OpenStudy (mertsj):

\[\frac{1}{6x^2}=\frac{1}{2x}+\frac{7}{6x^2}\]

OpenStudy (mertsj):

Do I have it right?

OpenStudy (anonymous):

yes.

OpenStudy (mertsj):

Common denominator is 6x^2. Let's multiply each term by that.

OpenStudy (mertsj):

\[6x^2\times\frac{1}{6x^2}=6x^2\times\frac{1}{2x}+6x^2\times\frac{7}{6x^2}\]

OpenStudy (mertsj):

\[1=3x+7\] \[-6=3x\] \[-2=x\]

OpenStudy (anonymous):

thanks so much, I figured I had the wrong common denominator.